OpenStudy (anonymous):

Show how to apply the order of operation rules as you simplify the following expression. |-5| – 45 ÷ 3

OpenStudy (freckles):

first of all what can you tell me the meaning of |-5| is

OpenStudy (freckles):

after you can say what that is then you will have subtraction and division to choose from division comes first then subtraction

OpenStudy (anonymous):

1 Simplify 453 to 15 |−5|−15 2 Simplify −10

OpenStudy (anonymous):

45 ÷ 3 =15 |-5|=5 -10 this is my answer but it was wrong. Idont know why.

OpenStudy (anonymous):

@Juando do you understand what is wrong with this answer?

OpenStudy (anonymous):

she said be thinking about the result of subtracting a number from a negative value.

OpenStudy (anonymous):

Anyone???

Directrix (directrix):

|-5| – 45 ÷ 3 ------------- @dwatts158 Your answer is wrong because you divided 45 by 3 and added it to |-5|. You should divide -45 by 3 and add it to |-5|

Directrix (directrix):

Think of subtraction as the addition of negative numbers.

Directrix (directrix):

So, this |-5| – 45 ÷ 3 is really |-5| + ( – 45) ÷ 3 which gives 5 - 15 = -10

Directrix (directrix):

Attached is your work with a sign correction. @dwatts158